OpenStudy (tywower):

A sample of gas has a volume of 20.0 liters at 22.0° C. If the pressure remains constant, what is the volume at 100.0° C?

OpenStudy (tywower):

15.8 liters 24.4 liters 25.3 liters 90.1 liters

OpenStudy (tywower):

@ganeshie8 !!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!1

OpenStudy (tywower):

Hii australopithecus!

OpenStudy (tywower):

@superhelp101

OpenStudy (australopithecus):

Use Ideal gas law. PV = nRT http://en.wikipedia.org/wiki/Ideal_gas_law You know pressure and moles of the substance is constant. The only two things changing are volume and temperature you know P1 = P2, also n1 = n2 Set up two ideal gas equations each for both different conditions, make them equal to terms P, n and R Since you know both formulas have the same pressure and moles you can set them both equal to each other (thus canceling out n1, P1, n2 and P2 and R), and solve for V2. Remember you know V1, you know T2, T1.

OpenStudy (australopithecus):

Gay-Lusacc's law relates changes of temperature and pressure, pressure is constant in this problem thus it does not relate to this specific problem

OpenStudy (australopithecus):

The right law to use is charlies law.

OpenStudy (anonymous):

I think the answer is C due to Charles law which is [v1/t1= v2/t2] so 100 * 20 = 200 then divide that by 22. Hope that helped. Not 100% sure though

OpenStudy (australopithecus):

better if you can derive these formulas for yourself using ideal gas law tbh

OpenStudy (anonymous):

@Australopithecus sorry I mixed up the laws, thanks for letting me know.

OpenStudy (anonymous):

I was always taught to memorize them, but ya deriving is good practice too.

OpenStudy (tywower):

UGHH idk who to give the medal to u both were so helpful!

OpenStudy (australopithecus):

\[\frac{P_1}{nR} =\frac{V_1}{T_1}\ and\ \frac{P_2}{nR} =\frac{V_2}{T_2}\] Since P1 and P2 are constants, moles dont change and R is a constant \[\frac{V_1}{T_1} = \frac{V_2}{T_2}\] Sub in the values you know and solve algebraically for V2 Using the following form (this case I am solving for x) \[\frac{x}{b} = \frac{c}{d}\] \[\frac{x}{b}*b = \frac{c}{d}*b\] \[\frac{x*b}{b} = \frac{c*b}{d}\] \[x*1 = x = \frac{c*b}{d}\]

OpenStudy (australopithecus):

I dont think I could be any more explicit in explaining this to you. Hope it helps